# Deterministic $(2/3-\varepsilon)$-Approximation of Matroid Intersection

### v1

@@ 4 $(2/3-\varepsilon)$-Approximation Algorithm using Nearly-Linear Independence-Oracle Queries
In this section, by providing a nearly-linear-independence-query $(2/3-\varepsilon)$-approximation algorithm for the matroid intersection problem, we prove Theorem 1, which we restate here.
See 1
We use the following lemma to prove Theorem 1.
Lemma 17.
@@
We argue by contraposition. Let $S\in\mathcal{I}_{1}\cap\mathcal{I}_{2}$ be a common independent set with $|S|<\frac{2}{3}r$. Then, we have $\frac{2|S|}{r-|S|}+2<6$. Thus, by Lemma 6, there exists an augmenting path of length at most $\frac{2|S|}{r-|S|}+2<6$ in $G(S)$. This completes the proof. ∎
Now, we give a proof of Theorem 1. See Algorithm 2 for the pseudocode of our algorithm.
Proof of Theorem 1.
@@
We give an algorithm to find a $(2/3-\varepsilon)$-approximate solution for the matroid intersection problem. In our algorithm, we first compute a maximal common independent set $S\in\mathcal{I}_{1}\cap\mathcal{I}_{2}$, which uses $O(n)$ independence oracle queries. Here, let $\bar{r}$ be the size of this maximal common independent set $S$. We note that $r\leq 2\bar{r}$ since any maximal common independent set is at least half the size of a maximum common independent set (see e.g., [KV06, Proposition 13.26]).
@@
Then, we compute the distance layers $D_{1}$, $D_{2}$, and $D_{3}$ in the exchange graph $G(S)$, which, by Lemma 8, uses $O(n+r\log r)$ independence oracle queries. Here, there is no augmenting path of length $2$ in $G(S)$, since $S$ is a maximal common independent set. Thus, shortest augmenting paths in $G(S)$ have length at least $4$.
@@
Next, by checking whether $S+v\in\mathcal{I}_{2}$ for each $v\in D_{3}$, we check whether the $(s,t)$-distance is $4$ or not, which uses $O(n)$ independence oracle queries. If the $(s,t)$-distance is more than $4$, then $S$ is already a $2/3$-approximate solution by Lemma 17, so we output $S$. Otherwise, we apply Lemma 15 in which we set the parameter $p=\varepsilon\bar{r}$, which uses $O(n/\varepsilon)$ independence oracle queries. [^8] We apply Refine (Algorithm 5) only $O(1/\varepsilon)$ times. Then, we obtain a partial augmenting set $\Phi=(B_{1},A_{1},B_{2})$. By applying Lemma 14 for the obtained partial augmenting set $\Phi$, using $O(n)$ independence oracle queries, we can find an augmenting set $\Pi=(B^{\prime}_{1},A^{\prime}_{1},B^{\prime}_{2})$ such that $B^{\prime}_{1}\subseteq B_{1}$, $A^{\prime}_{1}\subseteq A_{1}$ and $B^{\prime}_{2}=B_{2}$. Then, we output the set $S\oplus\Pi$.
@@
By Theorem 10, the set $S\oplus\Pi$ is a common independent set. Now, we show that the obtained solution $S\oplus\Pi$ is a $(2/3-\varepsilon)$-approximate solution.
@@
By the condition $(iii)$ in Lemma 15, there is a maximal augmenting set $\tilde{\Pi}=(\tilde{B_{1}},\tilde{A_{1}},\tilde{B_{2}})$ in $G(S)$ such that $|\tilde{B_{1}}|\leq|B_{1}|$. By Theorems 10 and 12 and Lemma 17, $S\oplus\tilde{\Pi}$ is a $2/3$-approximate solution, so we have $|S\oplus\tilde{\Pi}|\geq\frac{2}{3}r$. Moreover, by the condition $(ii)$ in Lemma 15, we have $|\tilde{B_{1}}|\leq|B_{1}|\leq|B_{2}|+p=|B^{\prime}_{2}|+p=|B^{\prime}_{1}|+p$. Thus, we have $|S\oplus\tilde{\Pi}|-|S\oplus\Pi|=|\tilde{B_{1}}|-|B^{\prime}_{1}|\leq p=\varepsilon\bar{r}\leq\varepsilon r$. Therefore, we have $|S\oplus\Pi|\geq(2/3-\varepsilon)r$, which completes the proof. ∎
@@
6 Apply Lemma 14 for the obtained partial augmenting set $\Phi=(B_{1},A_{1},B_{2})$ and then obtain an augmenting set $\Pi=(B^{\prime}_{1},A^{\prime}_{1},B^{\prime}_{2})$ such that $B^{\prime}_{1}\subseteq B_{1}$, $A^{\prime}_{1}\subseteq A_{1}$ and $B^{\prime}_{2}=B_{2}$.
@@ 5 $(2/3-\varepsilon)$-Approximation Algorithm using Linear Rank-Oracle Queries
In this section, by providing a linear-rank-query $(2/3-\varepsilon)$-approximation algorithm for the matroid intersection problem, we prove Theorem 2. We show that Algorithm 2 can be implemented as a $(2/3-\varepsilon)$-approximation algorithm with linear rank oracle queries. We restate Theorem 2 here.
See 2
Chakrabarty–Lee–Sidford–Singla–Wong [CLS+19] showed that a $(1-\varepsilon)$-approximate solution can be obtained with $O(\frac{n}{\varepsilon}\log n)$ rank oracle queries by using the binary search technique; see [CLS+19, Theorem 17]. In their $(1-\varepsilon)$-approximation algorithm, they find some edges using the binary search technique. For each edge, this requires $O(\log n)$ rank oracle queries. On the other hand, we show that a $(2/3-\varepsilon)$-approximate solution can be obtained with only $O(n/\varepsilon)$ rank oracle queries. In our $(2/3-\varepsilon)$-approximation algorithm, we find an almost maximal augmenting set without needing to identify any specific edges. Hence, the rank oracle query complexity of our algorithm is linear without any logarithmic factor.
We use the following lemma to prove Theorem 2.
Lemma 18.
@@
Proof.
Our algorithm for computing $D_{1},D_{2}$, and $D_{3}$ is almost the same as Algorithm 1. Since a single rank oracle query can determine whether a given subset is independent or not, Algorithm 1 except for computing $D_{2}$ can be implemented with $O(n)$ rank oracle queries. Now, we show how $D_{2}$ can be computed with $O(n)$ rank oracle queries. For each $u\in S$, we check whether $\textrm{rank}_{\mathcal{M}_{2}}(S+D_{1}-u)\geq\textrm{rank}_{\mathcal{M}_{2}}(S)$ holds or not. If $\textrm{rank}_{\mathcal{M}_{2}}(S+D_{1}-u)\geq\textrm{rank}_{\mathcal{M}_{2}}(S)$ holds, then there exists an element $v\in D_{1}$ such that $S+v-u\in\mathcal{I}_{2}$, and consequently $u\in D_{2}$. Otherwise, there does not exist an element $v\in D_{1}$ such that $S+v-u\in\mathcal{I}_{2}$, and consequently $u\notin D_{2}$. This completes the proof. ∎
Proof of Theorem 2.
Now, we show how Algorithm 2 can be implemented with $O(n/\varepsilon)$ rank oracle queries. As mentioned in the proof of Theorem 1, all parts of Algorithm 2, except for Line 2, can be implemented with $O(n/\varepsilon)$ independence oracle queries. Since a single rank oracle query can determine whether a given subset is independent or not, we can implement all parts of Algorithm 2, except for Line 2, with $O(n/\varepsilon)$ rank oracle queries. In addition, by Lemma 18, we can implement Line 2 with $O(n)$ rank oracle queries, which completes the proof. ∎
@@ 6 $(2/3-\varepsilon)$-Approximation Algorithm in the Semi-Streaming Model
In this section, by providing a $(2/3-\varepsilon)$-approximation algorithm for the matroid intersection problem in the semi-streaming model, we prove Theorem 3. We show that Algorithm 2 can also be implemented in the semi-streaming model as a $(2/3-\varepsilon)$-approximation algorithm using a constant number of passes. We restate Theorem 3 here.
See 3
Before we present how Algorithm 2 can be implemented in the semi-streaming model, we describe the outline of Blikstad’s [Bli21] algorithm of Lemma 15. This is because some of his results will be used in our argument later. In this paper, we skip the proof of the correctness of the algorithm; see [Bli21, Sections 3.1 and 3.3] for the proof. Recall that, in Lemma 15, we apply Refine only $O(1/\varepsilon)$ times. See Algorithm 5 for the pseudocode of Refine.
In Blikstad’s algorithm, we maintain three types of elements in each layer (see [Bli21, Section 3.1]):
@@ 4 $A_{k}\leftarrow A_{k}-A$, $R_{2k}\leftarrow R_{2k}+A$
Algorithm 3 RefineAB$(k)$ (from [Bli21, Algorithm 1]); called Refine1 in [CLS+19, Algorithm 9]
@@ 4 $A_{k}\leftarrow A_{k}+F_{2k}\setminus A$, $F_{2k}\leftarrow A$
Algorithm 4 RefineBA$(k)$ (from [Bli21, Algorithm 2]); called Refine2 in [CLS+19, Algorithm 10]
@@ 10       RefineAB$(k)$
Algorithm 5 Refine (from [Bli21, Algorithms 3 and 4])
Now, by providing how Algorithm 2 can be implemented in the semi-streaming model, we give a proof of Theorem 3.
Proof of Theorem 3.
@@
Even if we replace Lines 55 in Refine with UpdateABA$(k)$ (Algorithm 6), the procedure Refine can correctly find a partial augmenting set that satisfies the conditions $(i)$$(iii)$ in Lemma 15.
Proof.
Consider the order of the elements in $F_{2k+1}$ satisfying the following condition: an element added to $B_{k+1}$ in the first for loop of UpdateABA$(k)$ appears before any element that is not added to $B_{k+1}$ in that loop. Suppose that, in the execution of Lines 55 in Refine, the elements in $F_{2k+1}$ arrive in this order. Then, the changes of types of elements in $D_{2k+1}$ by UpdateABA$(k)$ call are exactly same as by the execution of Lines 55 in Refine (Algorithm 5). Regardless of the order of the elements in the for loop in Line 5, the implementation of Refine in Algorithm 5 can find a desired partial augmenting set. Therefore, Refine with Lines 55 replaced by UpdateABA$(k)$ correctly finds a desired partial augmenting set, which completes the proof. ∎
@@ 7             $R_{2k+1}\leftarrow R_{2k+1}+x$, $F_{2k+1}\leftarrow F_{2k+1}-x$
After applying Refine $O(1/\varepsilon)$ times, we obtain a partial augmenting set $\Phi=(B_{1},A_{1},B_{2})$ such that $|B_{1}|-|B_{2}|\leq\varepsilon\bar{r}$. Then, by applying Lemma 14 for the obtained partial augmenting set $\Phi=(B_{1},A_{1},B_{2})$, we obtain an augmenting set $\Pi=(B^{\prime}_{1},A^{\prime}_{1},B^{\prime}_{2})$ such that $B^{\prime}_{1}\subseteq B_{1}$, $A^{\prime}_{1}\subseteq A_{1}$ and $B_{2}=B^{\prime}_{2}$. The algorithm in Lemma 14 can be implemented without an additional pass of the stream, because it only accesses the set $S$ and the partial augmenting set $\Phi=(B_{1},A_{1},B_{2})$; see [CLS+19, Proof of Lemma 38].
Finally, we output the set $S\oplus\Pi$. By the same argument as Proof of Theorem 1, the set $S\oplus\Pi$ is a $(2/3-\varepsilon)$-approximate solution, which completes the proof. ∎
@@ 7 Concluding Remarks
We have observed that a $(2/3-\varepsilon)$-approximate solution can be obtained by terminating Blikstad’s [Bli21] algorithm early. Then, we obtained a deterministic nearly-linear-independence-query $(2/3-\varepsilon)$-approximation algorithm for the matroid intersection problem.
We were also able to use this observation in the streaming model of computation. Then, we obtain a $(2/3-\varepsilon)$-approximation constant-pass semi-streaming algorithm for the matroid intersection problem. This result is a generalization of the $(2/3-\varepsilon)$-approximation bipartite matching semi-streaming algorithm of Feigenbaum–Kannan–McGregor–Suri–Zhang [FKM+05], who initiated the study of matching algorithms in the semi-streaming model. Soon after, McGregor [McG05] first obtained a $(1-\varepsilon)$-approximation constant-pass semi-streaming algorithm for the maximum matching problem. Since then, there have been many studies for $(1-\varepsilon)$-approximation constant-pass semi-streaming algorithms for the maximum matching problem in the literature [EKS09, AG11, EKMS12, Kap13, Tir18, GKMS19, ALT21, FMU22, HS23]. Then, it is natural to ask whether we can obtain a $(1-\varepsilon)$-approximation constant-pass semi-streaming algorithm for the matroid intersection problem.

### v3

@@ 4 $(2/3-\varepsilon)$-Approximation Algorithm using Nearly-Linear Independence-Oracle Queries
In this section, by providing a nearly-linear-independence-query $(2/3-\varepsilon)$-approximation algorithm for the matroid intersection problem, we prove Theorem 1.1, which we restate here.
See 1.1
We use the following lemma to prove Theorem 1.1.
Lemma 4.1.
@@
We argue by contraposition. Let $S\in\mathcal{I}_{1}\cap\mathcal{I}_{2}$ be a common independent set with $|S|<\frac{2}{3}r$. Then, we have $\frac{2|S|}{r-|S|}+2<6$. Thus, by Lemma 2.3, there exists an augmenting path of length at most $\frac{2|S|}{r-|S|}+2<6$ in $G(S)$. This completes the proof.
Now, we give a proof of Theorem 1.1. See Algorithm 2 for the pseudocode of our algorithm.
Proof 4.3 (Proof of Theorem 1.1).
@@
We give an algorithm to find a $(2/3-\varepsilon)$-approximate solution for the matroid intersection problem. In our algorithm, we first compute a maximal common independent set $S\in\mathcal{I}_{1}\cap\mathcal{I}_{2}$, which uses $O(n)$ independence oracle queries. Here, let $\bar{r}$ be the size of this maximal common independent set $S$. We note that $r\leq 2\bar{r}$ since any maximal common independent set is at least half the size of a maximum common independent set (see e.g., [47, Proposition 13.26]).
@@
Then, we compute the distance layers $D_{1}$, $D_{2}$, and $D_{3}$ in the exchange graph $G(S)$, which, by Lemma 2.5, uses $O(n+r\log r)$ independence oracle queries. Here, there is no augmenting path of length $2$ in $G(S)$, since $S$ is a maximal common independent set. Thus, shortest augmenting paths in $G(S)$ have length at least $4$.
@@
Next, by checking whether $S+v\in\mathcal{I}_{2}$ for each $v\in D_{3}$, we check whether the $(s,t)$-distance is $4$ or not, which uses $O(n)$ independence oracle queries. If the $(s,t)$-distance is more than $4$, then $S$ is already a $2/3$-approximate solution by Lemma 4.1, so we output $S$. Otherwise, we apply Lemma 3.7 in which we set the parameter $p=\varepsilon\bar{r}$, which uses $O(n/\varepsilon)$ independence oracle queries. [^8] We apply Refine (Algorithm 5) only $O(1/\varepsilon)$ times. Then, we obtain a partial augmenting set $\Phi=(B_{1},A_{1},B_{2})$. By applying Lemma 3.6 for the obtained partial augmenting set $\Phi$, using $O(n)$ independence oracle queries, we can find an augmenting set $\Pi=(B^{\prime}_{1},A^{\prime}_{1},B^{\prime}_{2})$ such that $B^{\prime}_{1}\subseteq B_{1}$, $A^{\prime}_{1}\subseteq A_{1}$ and $B^{\prime}_{2}=B_{2}$. Then, we output the set $S\oplus\Pi$.
@@
By Theorem 3.2, the set $S\oplus\Pi$ is a common independent set. Now, we show that the obtained solution $S\oplus\Pi$ is a $(2/3-\varepsilon)$-approximate solution.
@@
By the condition $(iii)$ in Lemma 3.7, there is a maximal augmenting set $\tilde{\Pi}=(\tilde{B_{1}},\tilde{A_{1}},\tilde{B_{2}})$ in $G(S)$ such that $|\tilde{B_{1}}|\leq|B_{1}|$. By Theorems 3.2 and 3.4 and Lemma 4.1, $S\oplus\tilde{\Pi}$ is a $2/3$-approximate solution, so we have $|S\oplus\tilde{\Pi}|\geq\frac{2}{3}r$. Moreover, by the condition $(ii)$ in Lemma 3.7, we have $|\tilde{B_{1}}|\leq|B_{1}|\leq|B_{2}|+p=|B^{\prime}_{2}|+p=|B^{\prime}_{1}|+p$. Thus, we have $|S\oplus\tilde{\Pi}|-|S\oplus\Pi|=|\tilde{B_{1}}|-|B^{\prime}_{1}|\leq p=\varepsilon\bar{r}\leq\varepsilon r$. Therefore, we have $|S\oplus\Pi|\geq(2/3-\varepsilon)r$, which completes the proof.
@@
6 Apply Lemma 3.6 for the obtained partial augmenting set $\Phi=(B_{1},A_{1},B_{2})$ and then obtain an augmenting set $\Pi=(B^{\prime}_{1},A^{\prime}_{1},B^{\prime}_{2})$ such that $B^{\prime}_{1}\subseteq B_{1}$, $A^{\prime}_{1}\subseteq A_{1}$ and $B^{\prime}_{2}=B_{2}$.
@@ 5 $(2/3-\varepsilon)$-Approximation Algorithm using Linear Rank-Oracle Queries
In this section, by providing a linear-rank-query $(2/3-\varepsilon)$-approximation algorithm for the matroid intersection problem, we prove Theorem 1.2. We show that Algorithm 2 can be implemented as a $(2/3-\varepsilon)$-approximation algorithm with linear rank oracle queries. We restate Theorem 1.2 here.
See 1.2
Chakrabarty–Lee–Sidford–Singla–Wong [18] showed that a $(1-\varepsilon)$-approximate solution can be obtained with $O(\frac{n}{\varepsilon}\log n)$ rank oracle queries by using the binary search technique; see [18, Theorem 17]. In their $(1-\varepsilon)$-approximation algorithm, they find some edges using the binary search technique. For each edge, this requires $O(\log n)$ rank oracle queries. On the other hand, we show that a $(2/3-\varepsilon)$-approximate solution can be obtained with only $O(n/\varepsilon)$ rank oracle queries. In our $(2/3-\varepsilon)$-approximation algorithm, we find an almost maximal augmenting set without needing to identify any specific edges. Hence, the rank oracle query complexity of our algorithm is linear without any logarithmic factor.
We use the following lemma to prove Theorem 1.2.
Lemma 5.1.
@@
Proof 5.2.
Our algorithm for computing $D_{1},D_{2}$, and $D_{3}$ is almost the same as Algorithm 1. Since a single rank oracle query can determine whether a given subset is independent or not, Algorithm 1 except for computing $D_{2}$ can be implemented with $O(n)$ rank oracle queries. Now, we show how $D_{2}$ can be computed with $O(n)$ rank oracle queries. For each $u\in S$, we check whether $\textrm{rank}_{\mathcal{M}_{2}}(S+D_{1}-u)\geq\textrm{rank}_{\mathcal{M}_{2}}(S)$ holds or not. If $\textrm{rank}_{\mathcal{M}_{2}}(S+D_{1}-u)\geq\textrm{rank}_{\mathcal{M}_{2}}(S)$ holds, then there exists an element $v\in D_{1}$ such that $S+v-u\in\mathcal{I}_{2}$, and consequently $u\in D_{2}$. Otherwise, there does not exist an element $v\in D_{1}$ such that $S+v-u\in\mathcal{I}_{2}$, and consequently $u\notin D_{2}$. This completes the proof.
Proof 5.3 (Proof of Theorem 1.2).
Now, we show how Algorithm 2 can be implemented with $O(n/\varepsilon)$ rank oracle queries. As mentioned in the proof of Theorem 1.1, all parts of Algorithm 2, except for Line 2, can be implemented with $O(n/\varepsilon)$ independence oracle queries. Since a single rank oracle query can determine whether a given subset is independent or not, we can implement all parts of Algorithm 2, except for Line 2, with $O(n/\varepsilon)$ rank oracle queries. In addition, by Lemma 5.1, we can implement Line 2 with $O(n)$ rank oracle queries, which completes the proof.
@@ 6 $(2/3-\varepsilon)$-Approximation Algorithm in the Semi-Streaming Model
In this section, by providing a $(2/3-\varepsilon)$-approximation algorithm for the matroid intersection problem in the semi-streaming model, we prove Theorem 1.3. We show that Algorithm 2 can also be implemented in the semi-streaming model as a $(2/3-\varepsilon)$-approximation algorithm using a constant number of passes. We restate Theorem 1.3 here.
See 1.3
Before we present how Algorithm 2 can be implemented in the semi-streaming model, we describe the outline of Blikstad’s [11] algorithm of Lemma 3.7. This is because some of his results will be used in our argument later. In this paper, we skip the proof of the correctness of the algorithm; see [11, Sections 3.1 and 3.3] for the proof. Recall that, in Lemma 3.7, we apply Refine only $O(1/\varepsilon)$ times. See Algorithm 5 for the pseudocode of Refine.
In Blikstad’s algorithm, we maintain three types of elements in each layer (see [11, Section 3.1]):
@@ 4 $A_{k}\leftarrow A_{k}-A$, $R_{2k}\leftarrow R_{2k}+A$
Algorithm 3 RefineAB$(k)$ (from [11, Algorithm 1]); called Refine1 in [18, Algorithm 9]
@@ 4 $A_{k}\leftarrow A_{k}+F_{2k}\setminus A$, $F_{2k}\leftarrow A$
Algorithm 4 RefineBA$(k)$ (from [11, Algorithm 2]); called Refine2 in [18, Algorithm 10]
@@ 10       RefineAB$(k)$
Algorithm 5 Refine (from [11, Algorithms 3 and 4])
Now, by providing how Algorithm 2 can be implemented in the semi-streaming model, we give a proof of Theorem 1.3.
Proof 6.1 (Proof of Theorem 1.3).
@@
Even if we replace Lines 55 in Refine with UpdateABA$(k)$ (Algorithm 6), the procedure Refine can correctly find a partial augmenting set that satisfies the conditions $(i)$$(iii)$ in Lemma 3.7.
Proof 6.2.
[…]
@@ 7             $R_{2k+1}\leftarrow R_{2k+1}+x$, $F_{2k+1}\leftarrow F_{2k+1}-x$
After applying Refine $O(1/\varepsilon)$ times, we obtain a partial augmenting set $\Phi=(B_{1},A_{1},B_{2})$ such that $|B_{1}|-|B_{2}|\leq\varepsilon\bar{r}$. Then, by applying Lemma 3.6 for the obtained partial augmenting set $\Phi=(B_{1},A_{1},B_{2})$, we obtain an augmenting set $\Pi=(B^{\prime}_{1},A^{\prime}_{1},B^{\prime}_{2})$ such that $B^{\prime}_{1}\subseteq B_{1}$, $A^{\prime}_{1}\subseteq A_{1}$ and $B_{2}=B^{\prime}_{2}$. The algorithm in Lemma 3.6 can be implemented without an additional pass of the stream, because it only accesses the set $S$ and the partial augmenting set $\Phi=(B_{1},A_{1},B_{2})$; see [18, Proof of Lemma 38].
Finally, we output the set $S\oplus\Pi$. By the same argument as Proof of Theorem 1.1, the set $S\oplus\Pi$ is a $(2/3-\varepsilon)$-approximate solution, which completes the proof.
@@ 7 Concluding Remarks
We have observed that a $(2/3-\varepsilon)$-approximate solution can be obtained by terminating Blikstad’s [11] algorithm early. Then, we obtained a deterministic nearly-linear-independence-query $(2/3-\varepsilon)$-approximation algorithm for the matroid intersection problem.
We were also able to use this observation in the streaming model of computation. Then, we obtain a $(2/3-\varepsilon)$-approximation constant-pass semi-streaming algorithm for the matroid intersection problem. This result is a generalization of the $(2/3-\varepsilon)$-approximation bipartite matching semi-streaming algorithm of Feigenbaum–Kannan–McGregor–Suri–Zhang [27], who initiated the study of matching algorithms in the semi-streaming model. Soon after, McGregor [50] first obtained a $(1-\varepsilon)$-approximation constant-pass semi-streaming algorithm for the maximum matching problem. Since then, there have been many studies for $(1-\varepsilon)$-approximation constant-pass semi-streaming algorithms for the maximum matching problem in the literature [25, 1, 24, 42, 59, 31, 8, 30, 40]. Then, it is natural to ask whether we can obtain a $(1-\varepsilon)$-approximation constant-pass semi-streaming algorithm for the matroid intersection problem.